# Towards Robust and Cost-Efficient Knowledge Unlearning for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora. However, training LLMs on human-written text entails significant risk of privacy and copyright violations, which demands an efficient machine unlearning framework to remove knowledge of sensitive data without retraining the model from scratch. While Gradient Ascent (GA) is widely used for unlearning by reducing the likelihood of generating unwanted information, the unboundedness of increasing the cross-entropy loss causes not only unstable optimization, but also catastrophic forgetting of knowledge that needs to be retained. We also discover its joint application under low-rank adaptation results in significantly suboptimal computational cost vs. generative performance trade-offs. In light of this limitation, we propose two novel techniques for robust and cost-efficient unlearning on LLMs. We first design an Inverted Hinge loss that suppresses unwanted tokens by increasing the probability of the next most likely token, thereby retaining fluency and structure in language generation. We also propose to initialize low-rank adapter weights based on Fisher-weighted low-rank approximation, which induces faster unlearning and better knowledge retention by allowing model updates to be focused on parameters that are important in generating textual data we wish to remove. Through extensive experiments on the Training Data Extraction Challenge dataset using the GPT-Neo family of LLMs, we find that our approach can efficiently remove unwanted information with negligible deterioration in reasoning as well as generative capabilities.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have demonstrated substantial performance gains in downstream tasks as their model size and the amount of training data increase (Zhao et al., 2023). This has prompted many research groups to prioritize the collection of more data for pretraining them and the development of more larger models than before (Brown et al., 2020; Chowdhery et al., 2023; Rae et al., 2021; Smith et al., 2022). However, this approach has introduced significant privacy concerns, as these models tend to memorize data indiscriminately Carlini et al. (2021; 2022). For example, Personally Identifiable Information (e.g., names, phone numbers, and email addresses) can be easily extracted from the LLMs (Carlini et al., 2021). Additionally, OpenAI is currently facing several lawsuits related to alleged copyright infringement within their pretraining text corpus (Grynbaum & Mac, 2023). In response to these challenges and in light of the recent GDPR legislation on the right to be forgotten (Voigt & Von dem Bussche, 2017; Rosen, 2011; Villaronga et al., 2018), machine unlearning for LLMs has emerged as a critical and rapidly growing research field (Yao et al., 2023; Si et al., 2023a).
Exact unlearning can be achieved by filtering out sensitive information and retraining from scratch. However, this process is highly resource-intensive and inefficient, especially when data deletion requests occur serially rather than in large batches. As shown in Figure 1, this necessitates the development of approximate unlearning frameworks, where the goal is to remove knowledge of specific data instances from a pretrained model. Consequently, several methods for knowledge unlearning have been proposed (Wang et al., 2023; Chen & Yang, 2023; Jang et al., 2023). Among them, Jang et al. (2023) proposed a simple approach that fine-tunes LLMs using Gradient Ascent (GA) on the forgetting data, while also introducing several metrics to assess its effectiveness. They demonstrate that GA not only can be used for successful knowledge unlearning but also minimizes catastrophic forgetting of other knowledge within the LLMs. Nevertheless, GA still has significant drawbacks, such as the issue of training not converging due to unbounded loss and high unlearning cost caused by full fine-tuning for the entire LLM model.
@@
Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it lose knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion. After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other data points in the retain set ${\mathcal{D}}_{r}$ as well as preserve its previously acquired reasoning and generative capabilities.
@@
On the other hand, LoRA (Hu et al., 2021) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning for downstream tasks. The core idea behind LoRA is to freeze the pretrained LLM’s weights and train rank decomposition matrices within each layer during adaptation to downstream tasks. This approach not only significantly reduces the number of trainable parameters but also achieves performance comparable to full fine-tuning. Recently, Biderman et al. (2024) experimentally demonstrates that LoRA acts as a powerful regularization mechanism that prevents it from reaching the same performance levels as full fine-tuning. As a result, it results in less forgetting in retained knowledge. From this perspective, many of LoRA’s benefits are particularly advantageous in practical unlearning scenarios: 1) it significantly reduces the tuning cost for unlearning, and 2) it effectively prevents forgetting, making it a valuable method for achieving the important goal of retaining knowledge. Nonetheless, its application to LLM unlearning remains unexplored until now.
In this paper, we propose a novel approach for robust and efficient unlearning for LLMs by introducing a novel loss function and leveraging LoRA with a new initialization. First, we analyze the derivatives of the existing GA and highlight its shortcomings: 1) dispersed gradients causing unnecessary forgetting, and 2) an unbounded loss that fails to converge. To address these issues, we propose the Inverted Hinge loss (IHL). From the derivative analysis, we show that the IHL resolves all the problems of GA, enabling faster and robust unlearning. Second, we explore the intersection between LLM unlearning and cost-efficient fine-tuning using LoRA. We find that LoRA suffers from significantly suboptimal trade-offs between its downstream performance after unlearning and its ability to generate information requested for deletion. To achieve robust and efficient unlearning with LoRA, we introduce a new LoRA Initialization via Weighted Low-Rank Approximation. Extensive experiments on the LM-Eval and Dialogue benchmarks using Neo-GPT, Phi, and Llama2 LLMs show that our proposed Inverted Hinge Loss combined with LoRA achieves robust and cost-efficient unlearning than GA (See Figure 2).
@@
Figure 2: Compute cost for successful unlearning vs. post-unlearning downstream performances of GPT-Neo-125M. Each point represents a different forget set and LoRA rank (if used). Left: Accuracy averaged across 9 classification tasks (higher is better). Middle: F1 score averaged across 4 dialogue generation tasks (higher is better). Right: Perplexity on the validation set of the Pile dataset (lower is better). Dashed lines indicate the performances of the model prior to unlearning. Unlearning with vanilla LoRA leads to significant loss in performance compared to full-parameter unlearning. However, our proposed method using both the Inverted Hinge Loss and FLoRA performs competitively to full-unlearning in all three aspects while enjoying high parameter- and cost-efficiency of LoRA.
@@ 2 Related Work
Machine Unlearning. The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pre-trained model to discard information acquired from a specific subset of data. Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2021; Mehta et al., 2022; Chundawat et al., 2022; Cha et al., 2024). Recently, its significance has grown notably in Large Language Models (LLMs) due to its pivotal role in managing unintended memorization intrinsic to LLMs (Si et al., 2023b; Yao et al., 2024).
@@
First, Wang et al. (2023) introduces the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets. Chen & Yang (2023) proposes the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model frozen. While these approaches demonstrates impressive unlearning results across various datasets, they are hindered by significant drawbacks, such as the necessity to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) and the need to finetune models using them (Wang et al., 2023). Parameter merging-based methods recently garner attention as another category of unlearning for LLMs (Wu et al., 2023; Ilharco et al., 2022). Experimental results demonstrate that simple task arithmetic can facilitate cost-efficient unlearning for LLMs. However, these methods also exhibit limitations, such as achieving a relatively weak level of unlearning (Ilharco et al., 2022) or incurring higher computational costs to detect privacy neurons (Wu et al., 2023).
@@
In contrast, following an unlearning approach similar to that used in image classification (Golatkar et al., 2020; Cha et al., 2024), Jang et al. (2023) leverages Gradient Ascent (GA) to maximize the next-token prediction loss for a sequence of tokens from the forget data. Unlike in image classification, they demonstrate that LLMs do not suffer from forgetting of retained knowledge when using GA on the forget data. Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks.
@@
Parameter-Efficient Fine-Tuning. Fine-tuning pretrained LLMs to adapt towards specific downstream tasks and instructions comes at an intractable computational cost due to the large model size and complexity. To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023). Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2021; Zhang et al., 2023; YEH et al., 2023; Kopiczko et al., 2023; Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter. A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM. While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposes to initialize LoRA weights using the principal singular vectors and values of the linear weights. Inspired by this work, we propose to extract Fisher information from the data requested for unlearning and perform weighted low-rank approximation to prepare a LoRA initialization that induces faster unlearning and better preservation of general knowledge
@@ 3 Proposed Method
In this section, we introduce the notations and provide preliminaries on unlearning through Gradient Ascent (GA) and low-rank adaptation (LoRA). We then analyze the gradient properties of GA and highlight its inherent issues. Building on this analysis, we propose the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA. Finally, we present the idea of employing LoRA with a novel initialization strategy, thereby enabling cost-efficient unlearning.
@@ 3.1 Preliminaries
Problem and notation. Given a sequence of $T$ tokens ${\bm{x}}=(x_{1},x_{2},\dots,x_{T})$, a language model (LM) models the likelihood of the sequence via next token prediction: $p_{\theta}({\bm{x}})=\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{<t})$. After pretraining, we assume that an end-user has requested to delete a subset of the training set $\mathcal{D}_{f}\subset\mathcal{D}$, which we refer to as the forget set. Also, we consider a retain set $\mathcal{D}_{r}$ that represents the auxilary dataset which contains general knowledge the model should not forget after unlearning (e.g., Wikitext dataset (Merity et al., 2016)).
Gradient Ascent. Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, which leads to a simple yet effective baseline of gradient ascent (GA) (Cha et al., 2024; Jang et al., 2023). Unlike the usual gradient descent, GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
$\mathcal{L}_{GA}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t}))$
Note that in practice, the log-likelihood is implemented in terms of cross-entropy losses on incremental tokens. Thus, GA tunes the model towards maximizing the cross-entropy, the value of which is unbounded above by nature, making optimization ill-posed and unstable.
Low-Rank Adaptation. Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank, LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter. In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes
$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}$ (1)
During fine-tuning, the original weight ${\bm{W}}$ is kept frozen and only the low-rank factors ${\bm{A}}$ and ${\bm{B}}$ are updated via gradient descent. To ensure that the initial attachment of LoRA adapters does not alter the output of the LLM, LoRA defaults to initializing ${\bm{A}}$ with a Kaiming-uniform distribution (He et al., 2015) and ${\bm{B}}$ as the zero matrix. Note that after fine-tuning, LoRA adapters can simply be merged with the original weights ${\bm{W}}^{\prime}={\bm{W}}+{\bm{B}}{\bm{A}}$, thereby avoiding any additional latency during inference. In the context of unlearning, this feature of LoRA naturally promotes a layer of security as it prevents adversarial agents from tracking previous records of unlearning just by inspecting the model architecture.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated strong reasoning and memorization capabilities via pretraining on massive textual corpora. However, this poses risk of privacy and copyright violations, highlighting the need for efficient machine unlearning methods that remove sensitive data without retraining from scratch. While Gradient Ascent (GA) is commonly used to unlearn by reducing the likelihood of generating unwanted content, it leads to unstable optimization and catastrophic forgetting of retrained knowledge. We also find that combining GA with low-rank adaptation results in poor trade-offs between computational cost and generative performance. To address these challenges, we propose two novel techniques for robust and efficient unlearning for LLMs. First, we introduce Inverted Hinge loss, which suppresses unwanted tokens while maintaining fluency by boosting the probability of the next most likely token. Second, we develop a data-adaptive initialization for LoRA adapters via low-rank approximation weighted with relative Fisher information, thereby focusing updates on parameters critical for removing targeted knowledge. Experiments on the Training Data Extraction Challenge dataset using GPT-Neo models as well as on the TOFU benchmark with Phi-1.5B and Llama2-7B models demonstrate that our approach effectively removes sensitive information while maintaining reasoning and generative capabilities with minimal impact.
@@ 1 Introduction
Large Language Models (LLMs) exhibit substantial performance gains in downstream tasks with increasing model size and amount of pretraining data (Zhao et al., 2023). This has prompted extensive research on collecting high-quality textual corpora for LLM pretraining and developing larger models to an unprecedented scale (Brown et al., 2020; Chowdhery et al., 2023; Smith et al., 2022; Rae et al., 2021; Dubey et al., 2024). However, this approach has introduced significant privacy concerns due to LLMs’ tendency to memorize data indiscriminately (Carlini et al., 2021; 2023). For instance, Personally Identifiable Information (e.g., names, phone numbers, and email addresses) can be easily extracted from LLMs (Carlini et al., 2021). Additionally, OpenAI is facing multiple copyright infringement lawsuits due to unpermitted use of licensed articles during LLM pretraining (Grynbaum & Mac, 2023). In response to such challenges as well as increasing interest in one’s right to be forgotten (e.g., the GDPR legislation) (Voigt & Von dem Bussche, 2017; Rosen, 2011; Villaronga et al., 2018), machine unlearning for LLMs has emerged a critical and rapidly growing research field (Yao et al., 2023; Si et al., 2023).
One method for LLM unlearning would be to filter out sensitive data from the corpus and retrain the model from scratch, an approach known as exact unlearning. With unprecedentedly large models and pretraining datasets, this process is highly resource-intensive and can easily become intractable under the possibility of multiple data deletion requests made in a sequential manner. This motivates approximate unlearning, where the goal is to remove knowledge of specific data instances without retraining the model from scratch (Figure 1). In this regard, Jang et al. (2023) proposed a simple method that finetunes LLMs using Gradient Ascent (GA) on data requested for deletion, and also introduced $n$-gram-based metrics to assess its efficacy. However, the proposed GA method suffers significantly not only from unstable optimization due to the objective loss being unbounded, but also from high computational cost of full-finetuning all parameters within the LLM.
@@
Figure 1: The objective of LLM unlearning is to finetune a pretrained LLM such that it forgets knowledge on the forget set ${\mathcal{D}}_{f}$ consisted of textual data requested for deletion. After forgetting ${\mathcal{D}}_{f}$, the LLM must maintain knowledge of other relevant data points in the retain set ${\mathcal{D}}_{r}$ as well as reasoning and generative capabilities acquired from pretraining.
@@
Meanwhile, Low-Rank Adaptation (LoRA) has emerged as one of the most prominent techniques for parameter-efficient fine-tuning on downstream tasks (Hu et al., 2022). The core idea of LoRA is to freeze all pretrained weights and instead train low-rank decomposition matrices to model the weight changes in each linear layer, effectively reducing the number of trainable parameters and thus its memory cost. In addition to its efficiency, the low-rankness in LoRA also induces a powerful regularization (Biderman et al., 2024), which can be beneficial in the realm of LLM unlearning by stabilizing optimization and preventing catastrophic forgetting of other remaining knowledge. From this perspective, we conjecture that LoRA would be a valuable approach to consider in practical unlearning scenarios, yet its application to LLM unlearning remains unexplored.
In this paper, we explore LLM unlearning under the low-rank adaptation paradigm and propose two novel techniques for robust and efficient knowledge unlearning for LLMs. First, we analyze the derivatives of GA and highlight its shortcomings: 1) gradients increasing the probability of all other possible tokens cause unnecessary forgetting and 2) maximizing the next-token prediction loss involves unbounded optimization and can easily diverge. To address these issues, we propose the Inverted Hinge Loss (IHL) that aims to replace each token to unlearn with the next most-probable token, and show that IHL enables fast and stable tuning by resolving the issues of GA. Second, we find that the low-rank of LoRA imposes regularization too strong for LLM unlearning, and forcing the model to fully unlearn data points through extensive tuning leads to suboptimal cost vs. post-unlearning performance trade-offs. To accelerate unlearning in a controlled manner, we propose Fisher-Initialization of Low-rank Adapters (FILA), which data-adaptively assigns parameters responsible for generating unwanted information to adapters prior to tuning by decomposing the pretrained parameters weighted by the relative Fisher-information matrix. Our extensive experiments on unlearning samples in the Training Data Extraction Challenge dataset with GPT-Neo models as well as removing knowledge of fictitious author profiles using the TOFU benchmark with Phi-1.5B and Llama2-7B models consistently demonstrate that our proposed IHL combined with FILA significantly outperforms previous baselines in both efficiency and post-unlearning downstream performance.
@@
Figure 2: Compute cost for successful unlearning vs. post-unlearning downstream performances. We unlearn 32 randomly sampled sequences from the Training Data Extraction Challenge from GPT-Neo-125M. Each point represents a different forget set and LoRA rank (if used). Left: Accuracy averaged across 9 classification tasks (higher is better). Middle: F1 score averaged across 4 dialogue generation tasks (higher is better). Right: Perplexity on the validation set of the Pile dataset (lower is better). Dashed lines indicate the performances of the model prior to unlearning. Unlearning via gradient differences (GD) with vanilla LoRA leads to significant loss in performance compared to full-parameter GD unlearning due to lack of plasticity. However, our proposed method using both the Inverted Hinge Loss and Fisher-weighted LoRA initialization performs competitively to unlearning via full-finetuning in all three aspects while enjoying the cost-efficiency of LoRA.
@@ 2 Related Work
Machine Unlearning. The primary objective of machine unlearning (Cao & Yang, 2015) is to adapt a pretrained model to discard information acquired from a specific subset of data previously used for pretraining. Machine unlearning has garnered attention within neural network models dedicated to image classification (Golatkar et al., 2020; Tarun et al., 2023; Mehta et al., 2022; Chundawat et al., 2023; Cha et al., 2024). Recently, its significance has grown notably with Large Language Models (LLMs) due to crucial need for managing unintended memorization of pretraining data intrinsic to LLMs (Si et al., 2023; Yao et al., 2024b).
@@
Several machine unlearning algorithms have been proposed for LLMs, with most falling under the category of parameter optimization-based methods (Si et al., 2023). First, Wang et al. (2023) introduced the Knowledge Gap Alignment loss, which employs knowledge distillation between output predictions from models trained on different datasets. Chen & Yang (2023) proposed the unlearning layer and applied it to remove specific knowledge while keeping other parameters of the model intact. Liu et al. (2024) developed a two-stage framework: the harmful knowledge acquisition stage identifies and captures harmful knowledge within the model, and the knowledge negation stage removes that knowledge. While these approaches have demonstrated impressive unlearning results across various datasets, they are limited in applicability either by the need to retain entire training and forget datasets (Wang et al., 2023; Chen & Yang, 2023) or by the need to infer from a secondary model for knowledge distillation (Wang et al., 2023; Liu et al., 2024). Another category of LLM unlearning methods are tuning-free model-editing methods, an example of which leverages task arithmetic to suppress harmful text (Ilharco et al., 2023; Wu et al., 2023). While free from cost of tuning, these methods fails to show sufficient level of unlearning (Ilharco et al., 2023) or incur higher computational costs for detecting privacy neurons (Wu et al., 2023).
@@
In contrast, inspired by image classifier unlearning literature (Golatkar et al., 2020; Cha et al., 2024), Jang et al. (2023) adopted Gradient Ascent (GA) to unlearn from LLMs by maximizing the next-token prediction loss for all sequences in the forget set. Unlike in image classification, they demonstrate that LLMs show negligible forgetting of retained knowledge when using GA on the forget data. Consequently, they show that effective unlearning for LLMs is achievable solely using the forget dataset, while preserving general LLM knowledge and maintaining performance on downstream tasks. More recently, Yao et al. (2024a) demonstrated that incorporating GA with gradient descent on in-distribution data further enhances robustness against hyperparameters.
@@
Parameter-Efficient Fine-Tuning. Full-finetuning LLMs to adapt towards specific downstream tasks and instructions incurs an intractable computational cost due to the large model size and complexity. To alleviate this burden, previous work has developed Parameter-Efficient Fine-Tuning (PEFT) methods that adapt only a small portion of LLM parameters while freezing the other pretrained parameters intact (Liu et al., 2022b; Qiu et al., 2023; Liu et al., 2023). Inspired by the finding that fine-tuning LLMs exhibit a small intrinsic rank (Li et al., 2018; Aghajanyan et al., 2021), LoRA and many of its derivatives attach low-rank adapters to linear layers within the LLM (Hu et al., 2022; Zhang et al., 2023; Yeh et al., 2023; Kopiczko et al., 2024; yang Liu et al., 2024), with which the output from the original layer is linearly combined with that of the adapter. A key benefit of LoRA is that the adapter weights can be merged seamlessly to pretrained parameters after fine-tuning, such that the post-adaptation LLM shares the same inference cost as the base pretrained LLM. While most previous work have used random initialization of LoRA adapters, PiSSA (Meng et al., 2024) proposed to initialize LoRA weights using the principal singular vectors and values of the linear weights.
@@ 3 Proposed Method
In this section, we introduce useful notation and provide background information on unlearning with Gradient Ascent (GA) and Low-Rank Adaptation (LoRA). We then analyze derivatives of GA and highlight its inherent issues. Building upon this analysis, we present the Inverted Hinge loss and demonstrate its effectiveness in addressing the problems of GA. Lastly, we illustrate the method of initializing LoRA data-adaptively via Fisher-information, thereby enabling cost-efficient unlearning.
@@ 3.1 Preliminaries
Problem and notation. Given a sequence of $T$ tokens ${\bm{x}}=(x_{1},x_{2},\dots,x_{T})$, a language model (LM) models the likelihood of the sequence via next-token prediction: $p_{\theta}({\bm{x}})=\prod_{t=1}^{T}p_{\theta}(x_{t}|x_{<t})$. After pretraining, we assume that an end-user has requested to delete a subset of the training set $\mathcal{D}_{f}\subset\mathcal{D}$, which we refer to as the forget set. The retain set $\mathcal{D}_{r}$ refers to an auxiliary dataset that contains other relevant knowledge that must be retained after unlearning (e.g., Wikitext; Merity et al. 2017).
Gradient Ascent. Ideally, the LM must assign low probability to sequences in $\mathcal{D}_{f}$, leading to a simple yet effective baseline of Gradient Ascent (GA; Jang et al. 2023). GA unlearns a sequence of tokens ${\bm{x}}=(x_{1},\dots,x_{T})$ by maximizing the next-token prediction loss:
$\mathcal{L}_{\text{GA}}({\bm{x}})=-\sum_{t=1}^{T}\log(p_{\theta}(x_{t}|x_{<t})).$ (1)
In practice, the log-likelihood is computed using cross-entropy loss and thus we GA essentially minimizes the Negative Cross-Entropy (NCE) loss. Therefore, GA maximizing the next-token prediction loss involves unbounded optimization, leading to an ill-posed process with unstable tuning. While Gradient Difference (GD) aims to alleviate this instability by minimizing the next-token prediction loss for ${\mathcal{D}}_{r}$ alongside NCE on ${\mathcal{D}}_{f}$ as regularization, we find that the approach falls short of a fundamental solution, showing performance degradation as unlearning updates are made.
Low-Rank Adaptation. Based on the assumption that parameter changes due to LLM adaptation exhibits an intrinsic low-rank (Aghajanyan et al., 2021), LoRA models the change in parameters $\Delta{\bm{W}}\in\mathbb{R}^{d\times k}$ of each linear weight ${\bm{W}}\in\mathbb{R}^{d\times k}$ via a product of two low-rank matrices ${\bm{A}}\in\mathbb{R}^{r\times k}$ and ${\bm{B}}\in\mathbb{R}^{d\times r}$ where $r\ll\min(d,k)$ is the rank of the LoRA adapter. In other words, the output of the adapted linear layer given an input ${\bm{x}}$ becomes:
$({\bm{W}}+\Delta{\bm{W}}){\bm{x}}={\bm{W}}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}}.$
During fine-tuning, the original weight ${\bm{W}}$ is kept frozen and only the low-rank factors ${\bm{A}}$ and ${\bm{B}}$ are updated via gradient descent. To ensure that the initial attachment of LoRA adapters does not alter the output of the LLM, LoRA defaults to initializing ${\bm{A}}$ with a Kaiming-uniform distribution (He et al., 2015) and ${\bm{B}}$ as the zero matrix. After finetuning, LoRA adapters can simply be merged with the original weights ${\bm{W}}^{\prime}={\bm{W}}+{\bm{B}}{\bm{A}}$, thereby avoiding any additional latency during inference.